OpenStudy (anonymous):

I need help how do i do this problem? Solve the equation, using the domain of x = {-1, 0, 2}. 2x + 3y = 7

OpenStudy (anonymous):

Like if someone could explain to me on how to do this, please helpppp. :,(

OpenStudy (anonymous):

when x=-1 y=9/3=3 when x=0 y=7/3 when x=2 y=1